A passenger train leaves the train depot 2 hours after a freight train left the same depot. The freight train is traveling 20 mph slower than the passenger train. Find the rate of each train, if the passenger train overtakes the freight train in 3 hours.
Answer by lwsshak3(11628)

A passenger train leaves the train depot 2 hours after a freight train left the same depot. The freight train is traveling 20 mph slower than the passenger train. Find the rate of each train, if the passenger train overtakes the freight train in 3 hours.
**
let x=rate of speed of freight train
x+20=rate of speed of passenger train
distance=travel time*rate of speed (same for both trains)
..
5x=3(x+20)
5x=3x+60
2x=60
x=30
x+20=50
rate of speed of freight train=30 mph
rate of speed of passenger train=50 mph